40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 1 23:05:58 2011, kiwamu, Update, Green Locking, noise budget noise_budget.png
    Reply  Wed Apr 6 02:41:01 2011, kiwamu, Update, Green Locking, noise budget : some more noise  noise_budget.png
Message ID: 4482     Entry time: Fri Apr 1 23:05:58 2011     Reply to this: 4491
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: noise budget 

I made a coarse noise budget in order to decide our next actions for the X arm green locking.

So be careful, this is not an accurate noise budget !

 Some data are just coming from rough estimations and some data are not well calibrated.

noise_budget.png

 Assuming all the noise are not so terribly off from the true values, the noise at high frequency is limited by the dark noise of the PD or it already reaches to the IR inloop signal.

The noise at low frequency is dominated by the intensity noise from the transmitted green light although we thought it has been eliminated by the comparator.

In any case I will gradually make this noise budget more accurate  by collecting some data and calibrating them.

 

According to the plot what we should do are :

  * More accurate PD noise measurement

  * More accurate shot noise estimation

  * Searching for a cause of the small beat signal (see here) because a bigger beat signal lowers the PD noise.

  * Investigation of the Intensity noise

ELOG V3.1.3-